OpenStudy (anonymous):

derivative of (x^2)/(2+sqrtx)

OpenStudy (anonymous):

\[\frac{x ^{2}}{2+\sqrt{x}}\]\[\frac{(x ^{2})\prime(2+\sqrt{x})-(x ^{2})(2+\sqrt{x})\prime}{(2+\sqrt{x})^{2}}=\frac{2x(2+\sqrt{x})-(x ^{2})(\frac{1}{2\sqrt{x}})}{(2+\sqrt{x})^{2}}\]